OpenStudy (yrelhan4):

integration of 5tanx/(tanx-2) how do i approach this?

OpenStudy (abb0t):

\(5 \int\limits \frac{ \tan(x) }{ \tan(x)-2 }dx\) Is this correct?

OpenStudy (yrelhan4):

Yep.

OpenStudy (abb0t):

You COULD convert: \(tan(x)= \frac{ \sin(x) }{ \cos(x) }\) and try and simplify.

OpenStudy (yrelhan4):

Doesnt help i guess. Unless you can come up with something. sinx/(sinx-2cosx) dx is what i get.

OpenStudy (anonymous):

it is simple put tan(x)-2=u and simplify

OpenStudy (yrelhan4):

Ok so sec^2x dx=du what i get is (t+2)/(sec^2x*t) dt.. what next @irshadno1 ?

OpenStudy (anonymous):

1 + tan^2x = sec^2

OpenStudy (yrelhan4):

Yeah i got it. What we have to do is, put tanx=t, solve put sec^2=1+t^2... use partial fractions and simplify. I was forgetting to use partial fractions. You were the closest. Thank you.